OpenStudy (rahulmr):

for what values of 'a' will y=ax be a tangent to x^2+y^2+20x-10y+100=0 Can anyone help me with this question. Thanks.

OpenStudy (elonasushchik):

jeeZ...have any ideas?

OpenStudy (john_es):

Is this an x^3 or an x^2?

OpenStudy (rahulmr):

sorry about that. it was x^2

OpenStudy (baru):

substitute y=ax in x^2 + y^2....=0 you will get a quadratic equation in x

OpenStudy (baru):

since y=ax intersects x^2 + Y^2+...=0 at only one point, the discriminant should be zero.

OpenStudy (baru):

in other words, we should have repeated roots for the quadratic equation in x.

OpenStudy (john_es):

Good one. Do you know how to solve the problem with these hints?

OpenStudy (john_es):

You will have something like this, \[(1+a^2)x^2+(20-10a)x+100=0\]that you can solve with the quadratic equation

OpenStudy (john_es):

You will obtain, \[x=\frac{5 \left(a-2-\sqrt{-3 a^2-4 a}\right)}{a^2+1}\] And \[x=\frac{5 \left(a-2+\sqrt{-3 a^2-4 a}\right)}{a^2+1}\] But there must be only one point so, \[\sqrt{-3 a^2-4 a}=0\Rightarrow a=-4/3 ; a=0\]

OpenStudy (baru):

i graphed it, the answer is correct.. looks something like this